# Tumor-anchored deep feature random forests for out-of-distribution detection in lung cancer segmentation

### v1

@@ Abstract
Accurate segmentation of cancerous lesions from 3D computed tomography (CT) scans is essential for automated treatment planning and response assessment. However, even state-of-the-art models combining self-supervised learning (SSL) pretrained transformers with convolutional decoders are susceptible to out-of-distribution (OOD) inputs, generating confidently incorrect tumor segmentations, posing risks for safe clinical deployment. Existing logit-based methods suffer from task-specific model biases, while architectural enhancements to explicitly detect OOD increase parameters and computational costs. Hence, we introduce a plug-and-play and lightweight post-hoc random forests-based OOD detection framework called RF-Deep that leverages deep features with limited outlier exposure. RF-Deep enhances generalization to imaging variations by repurposing the hierarchical features from the pretrained-then-finetuned backbone encoder, providing task-relevant OOD detection by extracting the features from multiple regions of interest anchored to the predicted tumor segmentations. Hence, it scales to images of varying fields-of-view. We compared RF-Deep against existing OOD detection methods using 1,916 CT scans across near-OOD (pulmonary embolism, negative COVID-19) and far-OOD (kidney cancer, healthy pancreas) datasets. RF-Deep achieved AUROC >> 93.50 for the challenging near-OOD datasets and near-perfect detection (AUROC >> 99.00) for the far-OOD datasets, substantially outperforming logit-based and radiomics approaches. RF-Deep maintained similar performance consistency across networks of different depths and pretraining strategies, demonstrating its effectiveness as a lightweight, architecture-agnostic approach to enhance the reliability of tumor segmentation from CT volumes.
keywords:
Tumor segmentation, lung cancer, computed tomography, out-of-distribution detection, random forests, outlier exposure
journal: Medical Image Analysis
[…]
organization=Department of Medical Physics
Memorial Sloan Kettering Cancer Center,addressline=1275 York Avenue, city=New York City, postcode=10065, state=New York, country=USA
@@ 1 Introduction
Deep learning (DL) based medical image analysis is now applicable throughout the continuum starting from diagnosis, treatment planning, and response monitoring at follow-up to enhance care of patients with cancer. For instance, hybrid transformer-convolutional networks have shown to provide highly accurate tumor and healthy tissue segmentations from radiographic modalities (jiang2022self; Willemink2022; NguyenAAAI2023; YanWACV2023; Qayyum2023BHI; gu2024build). DL methods for medical applications are often developed using narrowly scoped datasets. Consequently, accuracy on in-distribution (ID) testing doesn’t translate to out-of-distribution (OOD) clinical cases impacted by different imaging acquisitions (Roschewitz2023Automatic; Koch2024), concept drifts (huang2022developing; sahiner2023data; gomez2024swintransformers), and label shifts like varying delineation guidelines (Nichyporuk2022MELBA; Lempart2023JACMP; Godau2025). Importantly, DL models also produce confidently inaccurate predictions (hendrycks2018deep; quinonero2022dataset; prabhu2022adapting_Nips2022; banerjee2023shortcuts), hindering ability of clinicians to leverage model uncertainties in utilizing AI predictions for clinical decision making (otles2021shift). Over-reliance on AI models by experts carries the risk of clinical expert deskilling (Budzyn2025Deskilling), exacerbating risk of harm to patients. These considerations underscore the critical need to provide reliably accurate methods that can contextualize task model predictions with respect to data instances used in prediction.
Segmentation accuracy metrics like Dice and Hausdorff distances, though useful to compare models’ performance on data cohorts, do not reflect well-calibrated uncertainty estimates even on ID datasets (Ren2024; Rangnekar2025SPIE). Model uncertainty visualization (goddard2012automation) has been shown to enhance consistency of clinical delineations (DeBiase2024). However, such measures can be unreliable in OOD scenarios, as demonstrated in three out of four OOD cases in Figure 1. Methods to explicitly flag OOD scans are often narrowly scoped to detect any variation from training data (Roschewitz2023Automatic; Godau2025); the model itself cannot generalize to common image variations not encountered in training, limiting their use for practical purposes. Our work overcomes the limitations of prior approaches by leveraging pretrained models to ensure sufficient generalization to common clinical imaging variations while enhancing robustness to far-OOD (disease in different anatomical sites) and near-OOD (different disease in same anatomic site) use cases. Our contributions are:
@@
Figure 1: Uncertainty maps for in-distribution (ID) and out-of-distribution (OOD) scans. (a) ID lung cancer cases (1–2) show concentrated boundary uncertainty; OOD cases (3–6: pulmonary embolism, COVID-19-negative, kidney cancer, healthy abdomen) exhibit concentrated, diffused, or misaligned patterns.
  • 1.
@@
    A post-hoc OOD detection approach using random forests with deep features (RF-Deep), trained with outlier exposure. Features are extracted from regions of interest (ROIs) anchored to the task-specific model’s predictions, enabling scan-level and task-relevant OOD filtering. Our approach reduced false detections, achieving AUROC >> 90 on challenging near-OOD datasets and near-perfect detection on the far-OOD datasets.
  • 2.
@@ 2 Related Works
Broadly, there are two classes of approaches for improving the generalizability of models for OOD detection. The first class consists of learner-based methods that leverage large corpus of training data, regularization techniques, and specialized network architectures to extract feature representations that are robust to data variations. The second class are filter-based methods that explicitly flag OOD cases at test time, preventing unreliable predictions and thereby improving overall reliability.
@@ 2.1 Learner-based methods
Learner-based methods focus on extracting robust feature representations from available training data such that the model can be applied across different domains without performance degradation. Domain-invariant feature representations can be extracted using: (a) specialized architectures (Saha2021EndToEndProstate; Simeth2023DominantIndexLesion), (b) data augmentation methods in image and feature space (Jung2019ProstateMixup; Chen2022EnhancingMRImageSegmentation; vaish2025dataagnostic), (c) regularized training with multi-task objectives (Panfilov2019RobustnessKneeMRI; Wen2024FromDenoisingTraining), as well as (d) SSL pretraining with numerous unlabeled examples to expose models to wider ranges of imaging and data variations than is practical for supervised training (hatamizadehSwinUNETR2022; haghighi2022dira; jiang2024self; Du2024SegVol).
Prior works on tumor segmentation have shown that use of enhanced network architectures combining attention mechanisms and residual connections that effectively increase network size produce robust generalization across different MR acquisitions (Saha2021EndToEndProstate; Simeth2023DominantIndexLesion). Another approach by Gu2021_DCANet_MICCAI explicitly trained a model to extract separate basis functions for various domains that were then combined using group convolutions to extract a generalizable representation applicable to unseen image domains for prostate gland segmentation.
An alternative approach to creating specialized architectures is to employ data augmentation and regularization tasks when training standard networks like the U-Net (Chen2022EnhancingMRImageSegmentation; Wen2024FromDenoisingTraining). Mixup is a type of data-agnostic augmentation strategy applicable either to the inputs, labels, as well as feature representations extracted from the input images (Jung2019ProstateMixup; vaish2025dataagnostic). Advanced feature augmentation strategies using normalizing flows extract feature statistics for creating augmented features in individual layers, enhancing domain generalization for prostate gland and eye fundus segmentation (SafRez_MixStyleFlow_MICCAI2025). Conceptually, the mixup and augmentation methods are general and could be combined with transformers to further enhance robustness to distribution shifts (Boone2023ROODMRI).
SSL pretraining is an altogether different approach compared to prior methods that employs multi-stage training to fully exploit the image information to extract distribution shift robust feature representations. This is accomplished either through pretraining with extensive data augmentations (Fedorov2021Tasting) or simply by exposing the model to a large corpus of unlabeled examples encompassing possible imaging variations (hatamizadehSwinUNETR2022; jiang2024self; gomez2024swintransformers). A different approach is a successive training strategy that explicitly addresses generalization to OOD by converting a portion of every encountered OOD data to ID data and successively refines the model to that dataset (Azizi2023REMEDIS). Unlike SSL, which aims to reduce need for large downstream datasets, this requires data for adapting to every new domain.
@@ 2.2 Filter-based OOD methods
Whereas learner-based methods seek to enhance models’ consistency across domains subject to distribution shifts, filter-based methods aim to enhance the confidence of predictions by filtering out data that differs from the training distribution. For this purpose, filter-based methods implement either a separate OOD model or a task-integrated OOD model to identify OOD data scans. Scans detected as OOD are removed, ensuring that models are applied only on ID-like data where predictions and uncertainties are valid, providing reliably safe performance required in clinical settings (berger2021confidence).
Similar to learner-based methods, filter-based methods have been applied to handle imaging acquisition differences (tardy2019uncertainty; gao2020response; nandy2021distributional). The vast majority of filter-based methods focused on disease classification from 2D images; method complexity varied from easy-to-compute Mahalanobis distances (tardy2019uncertainty), dedicated networks for detecting OOD cases (nandy2021distributional) to more complex instance-level retraining performed to assign metric distance to OOD scans by measuring response variations on the ID test set (gao2020response). A different approach involves using a single model to classify tasks on inlier data, while exposing the same model to outlier examples to detect OOD during inference (araujo2023few; roy2022does). The last approach is practical to implement when OOD cases are known as domain knowledge (e.g. optical tomography versus retinal fundus images) (araujo2023few), or when the outliers like infrequent disease conditions are deducible from the training as OOD (roy2022does).
Similar to classification tasks, Mahalanobis distance computed from feature embeddings has been used to detect OOD for lung lesion segmentation on CT scans of patients with COVID-19 (gonzalez2022distance). However, the high dimensionality of feature embeddings limits accuracy (roy2022does), due to difficulty in distinguishing subtle differences in appearance of different pathologies occurring in the same disease site (berger2021confidence). Dimensionality reduction methods including PCA (woodland2024dimensionality) as well as VQ-GAN (graham2023unsupervised; pinaya2022unsupervised) have been used to reduce the latent space and sharpen the separation of ID from OOD feature distributions. However, dimensionality reduction can suffer from high inductive bias (nalisnickMTGL19), reducing applicability to wider range of OOD scans. These dimensionality reduction approaches, including GANs, also require additional training with secondary datasets. Calibrated uncertainty estimation methods using computationally simple temperature scaling (karimi2022improving) as well as logit-based OOD scores (hendrycks17baseline; hendrycks2022logits) have been used for OOD detection. However, effectiveness of such methods are limited by inherent bias of the task-specific model use to estimate uncertainty (mehrtash2020confidence). Modified architectural formulations (yuan2023devil; Larrazabal2023MEEP) and generative diffusion methods (Nguyen2026) to detect OOD increase network parameters and require additional training. Finally, radiomic feature distance methods though useful for OOD detection at cohort level (konz2024radmetricmedicalimage; vasiliuk2023limitations) are not applicable to scan-level detections due to their difficulty in extracting subtle data variations while providing robust performance over common variations.
@@ 3 Framework overview and OOD task definition
Our goal for OOD detection was to develop a highly accurate scan-level OOD detection approach, which does not require large numbers of outlier samples, with generalizable applicability to images with different acquisitions and fields of view. Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the training dataset used to create the lung tumor segmentation model, and let $\mathcal{D}_{\text{out}}$ represent the scans that differ in pathology and anatomic sites. The objective is to determine whether a new scan $x$ belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$ at the scan level via a scoring mechanism provided by the random forest classifier.
Our approach combines the advantages of learner-based methods (by using the backbone of a segmentation model subject to SSL pretraining to extract imaging domain-agnostic features) with a filter-based classifier to reject conceptually different scans. An architecturally simple and lightweight random forest OOD classifier incorporates known outliers during training (hendrycks2018deep; thulasidasan2021effective; GuhaRoy2022) to enhance the classifier’s ability to flag unseen outliers. Scan-level OOD prediction for segmentation involves thousands of voxel-level predictions, which, while offering rich spatial context, also introduces memory constraints in processing and aggregating information from their large 3D volumes. Our approach addresses this issue by focusing the OOD detection in regions of interest (ROI) anchored to the predicted tumor segmentation, and then averaging the metric scores from those regions, which also allowed the approach to scale to images with varying field of views. Using regions anchored near the predicted tumors also allows the model to focus on differentiating features in relevant parts of the image, as opposed to distinguishing entire images, which in turn allows it to distinguish pathologies occurring in the same disease site.
@@ 3.1 Segmentation model
The segmentation framework employs a hybrid transformer-convolution architecture to leverage strengths of global contextual modeling with the transformer backbone as encoder and local spatial precision of the convolutional decoder. The backbone is a hierarchical Swin Transformer (liu2021swin) with depth configuration 221222-2-12-2 across four successive stages, with a patch size of $2\times 2\times 2$ voxels and a window size of $4\times 4\times 4$ voxels. Its hierarchical design progressively aggregates features at multiple spatial scales to capture fine-grained anatomical details. Windowed self-attention within each stage reduces complexity while preserving details and global context, critical for processing large volumetric images of size $128\times 128\times 128$ voxels. The backbone was pretrained with the self-distilled masked image transformer (SMIT) framework to enhance robustness to imaging and patient variations (jiang2022self). SMIT combines masked image prediction (he2021masked; xie2022simmim) with self-distillation (zhou2021ibot); described in Appendix A. The decoder, a 3D U-Net-based (ronneberger2015u) convolutional structure, was randomly initialized and fine-tuned with ID task-specific dataset. Following training, the entire model is frozen ([Uncaptioned image]) for subsequent segmentation and features-based OOD detection.

### v2

@@ Abstract
Accurate segmentation of cancerous lesions from 3D computed tomography (CT) scans is essential for automated treatment planning and response assessment. However, even state-of-the-art models combining self-supervised learning (SSL) pretrained transformers with convolutional decoders are susceptible to out-of-distribution (OOD) inputs, generating confidently incorrect tumor segmentations, posing risks to safe clinical deployment. Existing logit-based methods suffer from task-specific model biases, while architectural enhancements to explicitly detect OOD increase parameters and computational costs. Hence, we introduce a lightweight, plug-and-play post-hoc random forests-based OOD detection framework called RF-Deep that leverages deep features with limited outlier exposure. RF-Deep enhances generalization to imaging variations by repurposing the hierarchical features from the pretrained-then-finetuned backbone, providing task-relevant OOD detection by extracting the features from multiple regions of interest anchored to the predicted tumor segmentations. We compared RF-Deep against existing OOD detection methods using 2,056 CT scans across near-OOD (pulmonary embolism, negative COVID-19) and far-OOD (kidney cancer, healthy pancreas) datasets. RF-Deep achieved AUROC >> 93.50 for the challenging near-OOD datasets and near-perfect detection (AUROC >> 99.00) for the far-OOD datasets, substantially outperforming logit-based and radiomics approaches. RF-Deep maintained consistent performance across networks of different depths and pretraining strategies, demonstrating its effectiveness as a lightweight, architecture-agnostic approach to enhance the reliability of tumor segmentation from CT volumes.
@@ 1 Introduction
Deep learning (DL) based medical image analysis is now applicable throughout the continuum starting from diagnosis, treatment planning, and response monitoring at follow-up to enhance care of patients with cancer. For instance, hybrid transformer-convolutional networks have been shown to provide highly accurate tumor and healthy tissue segmentations from radiographic modalities (Jiang et al., 2022; Willemink et al., 2022; Nguyen et al., 2023; Yan et al., 2023; Qayyum et al., 2023; Gu et al., 2025). However, these methods are typically developed on narrowly-scoped datasets, and accuracy on in-distribution (ID) data does not translate to out-of-distribution (OOD) scenarios including different imaging acquisitions (Roschewitz et al., 2023; Koch et al., 2024), concept drifts (Huang et al., 2022; Sahiner et al., 2023; Gomez et al., 2025), and label shifts (Nichyporuk et al., 2022; Lempart et al., 2023; Godau et al., 2025). Importantly, DL models produce confidently inaccurate predictions (Hendrycks et al., 2019; Quiñonero-Candela et al., 2022; Prabhu et al., 2022; Banerjee et al., 2023), preventing clinicians from reliably using model uncertainties for decision making (Otles et al., 2021) and risking clinical expert deskilling through over-reliance on AI (Budzyń et al., 2025).
Common deployment failures that motivate our work include the Picture Archiving and Communication System (PACS) routing errors directing non-thoracic scans to lung-specific analysis pipelines and rare thoracic pathologies (e.g., pulmonary embolism) mimicking tumor appearances. Segmentation accuracy metrics like Dice and Hausdorff distances do not reflect well-calibrated uncertainty estimates, even on ID datasets (Ren et al., 2024; Rangnekar et al., 2025). While uncertainty visualizations can enhance clinical delineations (Goddard et al., 2012; Biase et al., 2024), such measures are unreliable in OOD scenarios, as shown in Figure 1. Existing OOD detection methods are often too sensitive, flagging any deviation from training data (Roschewitz et al., 2023; Godau et al., 2025), limiting their use for practical purposes.
Our work overcomes the limitations of prior approaches by leveraging pretrained models to ensure sufficient generalization to common clinical imaging variations while enhancing robustness to far-OOD (disease in different anatomical sites) and near-OOD (different disease in same anatomic site) use cases. We propose a tumor-anchored deep feature extraction framework for post-hoc OOD detection. Unlike prior logit-based or distance methods that operate on entire images or require architectural modifications, we extract hierarchical encoder features from multiple regions of interest anchored to predicted tumor segmentations. This enables task-relevant detection that distinguishes subtle pathological differences within the same anatomical site while scaling to images with varying fields-of-view. Our contributions are:
@@
Figure 1: Uncertainty maps for in-distribution (ID) and out-of-distribution (OOD) scans. (a) ID lung cancer scans (1–2) show concentrated boundary uncertainty; OOD scans (3–6: pulmonary embolism, COVID-19-negative, kidney cancer, healthy abdomen) exhibit concentrated, diffused, or misaligned patterns.
@@
  • A post-hoc OOD detection approach using random forests with deep features (RF-Deep), trained with outlier exposure. Features are extracted from regions of interest (ROIs) anchored to the task-specific model’s predictions, enabling scan-level and task-relevant OOD filtering. Our approach reduced false positive detections, achieving AUROC >> 90 on challenging near-OOD datasets and near-perfect detection on the far-OOD datasets.
@@ 2 Related Works
Approaches for OOD detection fall into two classes: learner-based methods that extract distribution-robust features through training strategies, and filter-based methods that explicitly flag OOD scans at test time.
@@ 2.1 Learner-based methods
Learner-based methods extract robust feature representations from available training data so models generalize across domains without performance degradation. Domain-invariant features can be obtained via (a) specialized architectures (Saha et al., 2021; Simeth et al., 2023), (b) data augmentation in image and feature space (Jung et al., 2019; Chen et al., 2022; Vaish et al., 2025), (c) regularized training with multi-task objectives (Panfilov et al., 2019; Wen et al., 2024), and (d) SSL pretraining on large unlabeled examples (Hatamizadeh et al., 2022; Haghighi et al., 2022; Jiang et al., 2024; Du et al., 2024).
For tumor segmentation, architectures combining attention mechanisms and residual connections have achieved robust generalization across MR acquisitions (Saha et al., 2021; Simeth et al., 2023). Gu et al. (2021) trained separate domain-specific basis functions combined via group convolutions for prostate gland segmentation. Alternatively, data augmentation (Safdari et al., 2025) and regularization (Jung et al., 2019; Vaish et al., 2025) have been applied to standard architectures like U-Net (Chen et al., 2022; Wen et al., 2024) to further enhance robustness to distribution shifts (Boone et al., 2023).
SSL pretraining takes a different approach, exploiting unlabeled data through extensive data augmentations (Fedorov et al., 2021) or exposure to diverse imaging variations (Hatamizadeh et al., 2022; Jiang et al., 2024; Gomez et al., 2025). Azizi et al. (2023) proposed successive refinement, converting OOD data to ID through iterative adaptation, although unlike SSL, this requires labeled data for each new domain.
@@ 2.2 Filter-based OOD methods
Filter-based methods reject OOD scans at test time, ensuring models operate only on ID-like data where predictions are reliable (Berger et al., 2021). These methods implement either separate OOD detectors or task-integrated detection, and have been applied to handle imaging acquisition differences (Tardy et al., 2019; Gao and Wu, 2020; Nandy et al., 2021). Most filter-based methods focused on 2D classification, with approaches ranging from Mahalanobis distances (Tardy et al., 2019) to dedicated OOD networks (Nandy et al., 2021) to instance-level retraining that measured response variations on ID test sets (Gao and Wu, 2020). A different approach involves using a single model to classify tasks on inlier data, while exposing the same model to outlier examples to detect OOD during inference (Roy et al., 2022; Araujo et al., 2023). This approach is practical when the OOD types are known from domain knowledge (e.g., optical tomography versus retinal fundus images) (Araujo et al., 2023), or when outliers such as infrequent disease conditions are identifiable from the training data (Roy et al., 2022).
For segmentation, Mahalanobis distance on feature embeddings has been applied to COVID-19 lung lesion detection (González et al., 2022), but the high dimensionality of feature embeddings limits accuracy, particularly for distinguishing pathologies within the same disease site (Berger et al., 2021; Roy et al., 2022). Dimensionality reduction via PCA (Woodland et al., 2024) or VQ-GAN (Pinaya et al., 2022; Graham et al., 2023) has been used to reduce the latent space and sharpen the separation of ID from OOD feature distributions, but introduces inductive bias (Nalisnick et al., 2019) and requires additional training. Logit-based scores (Hendrycks and Gimpel, 2017; Hendrycks et al., 2022) and temperature scaling (Karimi and Gholipour, 2022) offer computational simplicity but inherit task-model biases (Mehrtash et al., 2020). Architectural modifications (Yuan et al., 2023; Larrazabal et al., 2023) and generative diffusion methods (Nguyen et al., 2026) have improved OOD detection but increase network parameters and training complexity. Radiomic feature distances have shown utility for dataset-level OOD detection (Vasiliuk et al., 2023; Konz et al., 2024) but struggle with scan-level detections, as they cannot simultaneously capture subtle data variations while providing robust performance to common imaging variations.
@@ 3 Framework overview and OOD task definition
To address these limitations, our goal was to develop a highly accurate scan-level OOD detection approach, which does not require large numbers of outlier samples, with generalizable applicability to images with different acquisitions and fields of view. Let $\mathcal{D}_{\text{in}}$ denote the distribution of lung cancer CT scans, closely matching the training dataset used to create the lung tumor segmentation model, and let $\mathcal{D}_{\text{out}}$ represent the scans that differ in pathology and anatomic sites. Given a new scan $x$, the objective is to determine whether it belongs to $\mathcal{D}_{\text{in}}$ or $\mathcal{D}_{\text{out}}$ at the scan level using a scoring mechanism provided by a random forest (RF) classifier.
Our approach combines the strengths of learner-based and filter-based OOD detection methods. Specifically, we leverage the backbone of a segmentation model pretrained via self-supervised learning to extract imaging features that are largely domain-agnostic, and pair these features with a lightweight random forest classifier to reject conceptually dissimilar scans. The RF-based OOD classifier incorporates known outliers during training, an outlier exposure strategy shown to be effective in prior work (Hendrycks et al., 2019; Thulasidasan et al., 2021; Guha Roy et al., 2022). We adopt outlier exposure for three key reasons. First, in clinical deployment, common OOD failure modes (e.g., incorrect body region or scanner-related differences) are often predictable based on domain knowledge. Second, purely unsupervised methods (such as Mahalanobis distance–based approaches) frequently struggle with near-OOD scans (Table 1), where pathologies share anatomical context with the in-distribution data. Third, our method requires only a small number of outlier examples (as few as 20 scans) to achieve strong performance (Section 5.5), making outlier curation feasible in practice.
Importantly, we evaluate generalization to unseen OOD categories via leave-one-dataset-out experiments (Section 5.4.2), demonstrating that RF-Deep extends beyond the specific outliers encountered during training. Scan-level OOD prediction for segmentation involves thousands of voxel-level predictions, which, while offering rich spatial context, also introduces memory constraints in processing and aggregating information from their large 3D volumes. Our approach addresses this issue by focusing the OOD detection in regions of interest (ROI) anchored to the predicted tumor segmentation, and then averaging the metric scores from those regions, which also allowed the approach to scale to images with varying fields of view. Using regions anchored near the predicted tumors also allows the model to focus on differentiating features in relevant parts of the image, as opposed to distinguishing entire images, enabling distinction of pathologies occurring in the same disease site.
A key design consideration is RF-Deep’s dependence on the segmentation model’s predictions. We observed that on OOD scans, the segmentation model typically produced confidently incorrect tumor-like regions, rather than failing to produce any output. These predictions served as anchors for ROI extraction, and the extracted features from these regions are precisely what RF-Deep uses to identify the scan as OOD. In rare cases where the model produces no foreground predictions, the absence itself can serve as an implicit OOD signal, as ID scans with confirmed tumors would be expected to yield predictions. Across our evaluation datasets, the segmentation model produced at least one connected component on almost 90% of OOD scans, confirming that the tumor-anchored approach remains applicable even under distribution shift.
@@ 3.1 Segmentation model
The segmentation framework employs a hybrid transformer-convolution architecture to leverage strengths of global contextual modeling with the transformer backbone as encoder and local spatial precision of the convolutional decoder. The backbone is a hierarchical Swin Transformer (Liu et al., 2021) with depth configuration 221222-2-12-2 across four successive stages, with a patch size of $2\times 2\times 2$ voxels and a window size of $4\times 4\times 4$ voxels. Its hierarchical design progressively aggregates features at multiple spatial scales to capture fine-grained anatomical details. Windowed self-attention within each stage reduces complexity while preserving details and global context, critical for processing large volumetric images of size $128\times 128\times 128$ voxels. The backbone was pretrained with the self-distilled masked image transformer (SMIT) framework to enhance robustness to imaging and patient variations (Jiang et al., 2022). SMIT combines masked image prediction (He et al., 2021; Xie et al., 2022) with self-distillation (Zhou et al., 2022); described in Appendix A. The decoder, a 3D U-Net-based (Ronneberger et al., 2015) convolutional structure, was randomly initialized and fine-tuned on the ID task-specific dataset. Following training, the entire model is frozen ([Uncaptioned image]) for subsequent segmentation and feature-based OOD detection.